Put the equation in Standard Form with a positive x coefficient.
y+2=5/3(x+6)

Answer by stanbon(75887)

y+2=5/3(x+6)
y = (5/3)x + 10 - 2
----
y = (5/3)x + 8
-------------
Cheers,
Stan H.
------------